OpenStudy (anonymous):

IF \[A+B+C=\pi\] then prove that- \[cosA+cosB+cosC=1+4\sin \frac {A}{2}\sin \frac{B}{2} \sin \frac{C}{2}\]

OpenStudy (anonymous):

\[\cos A+\cos B+\cos C=\cos A+\cos B+\cos (\pi-A-B)\\=\cos A+\cos B-\cos (A+B)=\cos A+\cos B-\cos A \cos B+\sin A \sin B\\=1-(1-\cos A)(1-\cos B)+\sin A \sin B \\=1-4 \ \sin^2 \frac{A}{2} \sin^2 \frac{B}{2}+4\sin \frac{A}{2} \cos \frac{A}{2} \sin \frac{B}{2} \cos \frac{B}{2}\] now just factor out \[ 4 \ \sin \frac{A}{2} \sin \frac{B}{2}\] see what happens

OpenStudy (anonymous):

continuing the work of @mukushla one gets \[ 1-4 \ \sin^2 \frac{A}{2} \sin^2 \frac{B}{2}+4\sin \frac{A}{2} \cos \frac{A}{2} \sin \frac{B}{2} \cos \frac{B}{2}\\ 1+4 \ \sin \frac{A}{2} \sin \frac{B}{2}\left(-\sin \frac{A}{2} \sin \frac{B}{2}+ \cos \frac{A}{2} \cos \frac{B}{2} \right)\\ 1+4 \ \sin \frac{A}{2} \sin \frac{B}{2} \cos \left(\frac A 2 + \frac B 2\right) \\ 1+4 \ \sin \frac{A}{2} \sin \frac{B}{2} \cos \left(\frac \pi 2-\frac C 2\right) \\ 1+4 \ \sin \frac{A}{2} \sin \frac{B}{2} \sin\frac C 2 \\ \]